A bullet moving with a speed of 200m/s strikes abag of sand & remains in it. the masses of bag& bullet are 4.5kg&500g respectively. If the sand bag is free to move its speed is
1
Expert's answer
2012-09-13T10:57:24-0400
Let I be the momentum of the bullet and J be the momentum of a bag striked with it, M and m be mass of the bag and the bullet respectively, V and W velocities of the bullet and the striked bag. Then:

I = mV
J = (M+m)W
I = J
mV = (M+m)W
W = mV/(M+m)
W = 0.5*200/(4.5+0.5) = 20 m/s.

So, the bag's speed is 20 m/s.